# On the variational dual formulation of the Nash system and an adaptive convex gradient-flow approach to nonlinear PDEs

### v1

@@ 4. A toy example of the gradient flow performance
Here $c$ is a scalar parameter (in what follows, we tacitly assume $c\neq-\frac{1}{2}$; the special case $c=-\frac{1}{2}$ is addressed in Remark 5). It is obvious that (4.1) has two solutions:
@@
At time $\tilde{\mathsf{s}}^{*}_{1}=2$ the trajectory reaches $d=\frac{1}{2}$ at the boundary of the “DtP zone”, but at any $\mathsf{s}^{*}_{1}<2$ the base state produced by the DtP map is $\bar{U}^{2}=(0,0)$, cf. (4.18), i.e., the new base state coincides with the old one. However, a minor perturbation of the zero base state creates a bias and resolves the issue, allowing the scheme to converge to one of the solutions, as it did above [^5] The same issue was discussed in [21, Sec. 2.2.2] from a different perspective (not involving a gradient flow) for the dual problem in the simple example of the intersection of a circle with a vertical line.. This indicates that, in potential applications of the scheme to PDEs, certain base states should be avoided (e.g., to prevent excessive symmetry).
@@ 5. A brief description of the gradient flow scheme for the noise-free Nash system
$\int_{0}^{T}\left(\partial_{s}a^{1},\delta a\right)\,dt\\ =\int_{0}^{T}\left[(v^{1}-v_{0},\delta E)+(v^{1}\otimes v^{1},\delta B)\right]\,dt,\quad a^{1}(T,x,\mathsf{s})=0,\quad a^{1}(t,x,0)=0,$ (5.7)
@@
$v^{1}(t,x,\mathsf{s}):=(I+2B^{1}(t,x,\mathsf{s}))^{-1}(\bar{v}^{1}(t,x)-E^{1}(t,x,\mathsf{s})).$
@@
would allegedly be a solution to the Nash system (2.2). In the second case, we generate an infinite sequence of base states $\{\bar{v}^{k}\}$. If this sequence has accumulation points in some reasonable topology (say, in the weak $L^{2}$-topology), those points can be referred to as “generalized” solutions to the Nash system.
@@ Acknowledgement
This work originated from discussions held at Carnegie Mellon University while DV was visiting AA; DV acknowledges the hospitality of CMU. DV’s work was supported by CMUC [^6] https://doi.org/10.54499/UID/00324/2025 under FCT, grants UID/00324/2025 and UID/PRR/00324/2025, and by KAUST Research Funding (KRF), award ORFS-2024-CRG12-6430.3.

### v2

@@ 4. A toy example of the gradient flow performance
Here $c$ is a scalar parameter (in what follows, we tacitly assume $c\neq-\frac{1}{2}$; the special case $c=-\frac{1}{2}$ is addressed in Remark 4.1). It is obvious that (4.1) has two solutions:
@@
At time $\tilde{\mathsf{s}}^{*}_{1}=2$ the trajectory reaches $d=\frac{1}{2}$ at the boundary of the “DtP zone”, but at any $\mathsf{s}^{*}_{1}<2$ the base state produced by the DtP map is $\bar{U}^{2}=(0,0)$, cf. (4.18), i.e., the new base state coincides with the old one. However, a minor perturbation of the zero base state creates a bias and resolves the issue, allowing the scheme to converge to one of the solutions, as it did above [^8] The same issue was discussed in [22, Sec. 2.2.2] from a different perspective (not involving a gradient flow) for the dual problem in the simple example of the intersection of a circle with a vertical line.. This indicates that, in potential applications of the scheme to PDEs, certain base states should be avoided (e.g., to prevent excessive symmetry).
@@ 5. A brief description of the gradient flow scheme for the noise-free Nash system
$\int_{0}^{T}\left(\partial_{\mathsf{s}}a^{1},\delta a\right)\,dt\\ =\int_{0}^{T}\left[(v^{1}-v_{0},\delta E)+(v^{1}\otimes v^{1},\delta B)\right]\,dt,\quad a^{1}(T,x,\mathsf{s})=0,\quad a^{1}(t,x,0)=0,$ (5.7)
@@
$v^{1}(t,x,\mathsf{s}):=(I+2\mathfrak{U}^{*}\operatorname{div}a^{1}(t,x,\mathsf{s}))^{-1}(\bar{v}^{1}(t,x)-\partial_{t}a^{1}(t,x,\mathsf{s})).$ (5.8)
Here the variation $\delta a(t,x)$ should satisfy
$\delta a(T,x)=0.$
If we test (5.7) with compactly supported $\delta a$ and integrate by parts, we formally get the following nonlinear degenerate parabolic PDE system of $n$ equations:
$\partial_{\mathsf{s}}a^{1}=-\partial_{t}v^{1}-\nabla\mathfrak{U}(v^{1}\otimes v^{1}),$ (5.9)
with $v^{1}$ defined by (5.8). Testing now with arbitrary $\delta a$, we infer that the boundary (and initial) conditions are
$v^{1}(0,x,\mathsf{s})=v_{0}(x),\quad a^{1}(T,x,\mathsf{s})=0,\quad a^{1}(t,x,0)=0.$ (5.10)
@@
would allegedly be a solution to the Nash system (2.2). In the second case, we generate an infinite sequence of base states
$\bar{v}^{k}(t,x):=v^{k-1}(t,x,{\mathsf{s}}^{*}_{k-1}).$
If this sequence has accumulation points in some reasonable topology (say, in the weak $L^{2}$-topology), those points can be referred to as “generalized” solutions to the Nash system.
@@ Acknowledgement
This work originated from discussions held at Carnegie Mellon University while DV was visiting AA; DV acknowledges the hospitality of CMU. DV’s work was supported by CMUC [^9] https://doi.org/10.54499/UID/00324/2025 under FCT, grants UID/00324/2025 and UID/PRR/00324/2025, and by KAUST Research Funding (KRF), award ORFS-2024-CRG12-6430.3. AA thanks Marshall Slemrod for suggesting the Nash system for study.